A pie shaped(triangular) lake-front lot has a perimeter of 1800 feet. One side is 200 feet longer than the shortest side, while the third side is 400 feet longer than the shortest side. What are the lengths of all three sides?
Help would be much appreciated. =]

Answer by vertciel(183)

Hello there,
I'm just wondering: are there any actual triangular lake front lots out there? It'd be cool to see.
Anyways, let's denote x as the shortest side.
Shortest side = x
One side = 200 + x
Third Side = 400 + x
Perimeter of a Triangle = Sum of All sides
1800 = x + 200 + x + 400 + x
So all you need to do now is to solve for x, and than based on x, add the appriopriate integer to get the different sides.
Write back if you need help. Good luck!